OpenStudy (anonymous):

Julio takes his brother to a music store to buy cds. Together they have $50 to spend on cds. The music store sells CDs for $15 each. Write a equation for the amount of money, y, Julio and his brother will have left over after purchasing x amount of CDs.

OpenStudy (anonymous):

ok if they are selling cds for 15$ a piece then you have to divide 15 into 50 to see how many you are going to have. can you do that for me?

OpenStudy (anonymous):

i think it goes in 4 times right?

OpenStudy (anonymous):

oh wait 3

OpenStudy (anonymous):

gooood, that's better

OpenStudy (anonymous):

ok so we know that our money will only go so far as 3*15. we will make 15 the x so we can buy 3x =y cds. how much will you have left over after 3x?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

very good! so your equation is y=3x and your leftover is 5. good job!

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

who hold on

OpenStudy (anonymous):

i read that question wrong. this is qwriting an equation for hte amount of money you will have left over?

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

the anwsers on my paper look ike this; y=15x + 50

OpenStudy (anonymous):

ok i'm so sorry...it would be 50 - 3x = y.....the money you start with minus the cost of cd is what you have left over.

OpenStudy (anonymous):

hmmm.....that would mean x would have to = -3 to have 5 $ left.....if that's what it says on yoru paper i would go with it.....but that just doesn't seem right to me

OpenStudy (anonymous):

A. y=50-15 B. Y=15x +50 C.y= 50x +15 D. y =-15x it could be any of these

OpenStudy (anonymous):

the way i read this it would be y = 50-15x........ but of course that isn't in your list......the only one that would make sense would be B because D doesn't tell how much your have left plus it will be negative.

OpenStudy (anonymous):

sorry A. is supposed to have a x at the end of the 15, so does that make A. the answer?

OpenStudy (anonymous):

yep!!! y = 50-15x

OpenStudy (anonymous):

awesome thank you